OpenStudy (anonymous):

how to test symmetric over the line pi/2 with the equation r=6/(1+sintheta)

OpenStudy (anonymous):

\[r=6\sin \theta\] Multiply both sides by r \[r^2 = 6 r \sin \theta\] transform to rectangular coordinates. \[x^2+y^2 = 6y\] \[x^2+y^2-6y+9=9\] \[x^2+(y-3)^2=9\] A circle with center (0,3) and radius 3 is symmetric about the line x=0 (y-axis)

OpenStudy (anonymous):

wait, how did you get 6 sin \[\theta\] when the equation is \[6 \div 1 +\sin \theta\] and they want to test algebraically

OpenStudy (anonymous):

So, YOU CAN do TAT!

OpenStudy (anonymous):

i use the method of replacing \[(r,\theta) \to (-r,-\theta) \] the i get the final answer but it doesn't seem right because it is \[r=6\div1-\sin \theta \] it isn't the same, so I am confused even though I know it must be symmetric. because the graph is a parabola

OpenStudy (anonymous):

@dpaInc a medal

OpenStudy (anonymous):

@Rohangrr , she's right... the equation is